helpProblem 5-1 Future Value (LG5-1) 25 Compute the future value in year 7 of a...

Problem 5-1 Future Value (LG5-1) 

Compute the future value in year 7 of a $2,200 deposit in year 1, and another $1700 deposit at the end of year 4 using a 8 percent interest rate (Do not round intermediate calculations and round your final answer to 2 decimal places.) 


Problem 4 and 5-6 Present Value and Annuity Payments 

A local furniture store is advertising a deal in which you buy a $4,200 living room set with three years before you need to make any payments (no interest cost is incurred). 

How much money would you have to deposit now in a savings account earning 4 percent APR, compounded monthly, to pay the $4,200 bill in three years? (Do not round intermediate calculations and round your final answer to 2 decimal places.) 

How much would you have to deposit in the savings account each month to be able to pay the bill? (Do not round intermediate calculations and round your final answer to 2 decimal places.) 


19 Problem 5-33 Annuity Interest Rate (LG5-8) 

What's the interest rate of a 5 year, annual $5,300 annuity with present value of $21,5007 (Use a time value of money calculator or a spreadsheet. Round your answer to 2 decimal places.) 


Problem 5-17 Future Value (LG5-1) 

Given a 5 percent interest rate, compute the year 6 future value of deposits made in years 1, 2, 3, and 4 of $1,650, $1,850, $1,850, and $2,150, respectively. (Do not round intermediate calculations and round your final answer to 2 decimal places.)  
Answer #1

1.
=2200*1.08^6+1700*1.08^3=5632.6339104768

2.
=4200/(1+4%/12)^(12*3)=3725.80927009836

3.
=4200*(4%/12)/((1+4%/12)^(12*3)-1)=110.00073702874

4.
=RATE(5,-5300,21500)=7.40047825679219%

5.
=1650*1.05^5+1850*1.05^4+1850*1.05^3+2150*1.05^2=8866.532390625